Two gaseous streams enter a combining tube and leave as a single mixture. These data apply at the entrance section:


For one gas : A1 = 55 cm2 U1 = 400fps    v1 = 10 ft3/lbm

For another gas: A2 = 50 cm2  m2 = 12.22lbm/s    р = 00.12lbm/ft3

At exit section: U3 = 350 fps      v3 = 8ft3/lbm

Find (a) U2 at section 2 (b) mass flow rate at the exit


1
Expert's answer
2021-09-17T15:32:43-0400

"A_1 = 55 \;cm^2 = 0.0592 \;ft^2 \\

U_1 = 400 \;fps \\

v_1 = 10 \;ft^3/lb \\

P_1 = 0.1 \;\frac{lb}{ft^3} \\

A_2 = 50 \;cm^2 = 0.0538 \;ft^2 \\

m_2 = 12.22 lb/s \\

P_2 = 0.12 \;lb/ft^3 \\

U_3 = 350 \;fps \\

v_3 = 8 \;ft^3/lb \\

P_3 = \frac{1}{8} \; lb/ft^3"

(a) To find speed at section 2

Mass flow rate at section 2:

"m_2 = P_2A_2U_2 \\

U_2 = \frac{m_2}{P_2A_2} \\

= \frac{12.22 \;lb/ft^3}{0.12 \; lg/ft^3 \times 0.0538 \;ft^2} \\

= 1892.21 \;fps"

(b) To find mass flow rate at the exit

"m_1=P_1A_1U_1 \\

m_1 = 0.1 \times 0.0592 \times 400 = 2.368 \; lb/s"

Conservation of mass

"m_3 = m_1+m_2 \\

= 2.368 + 12.22 \\

= 14.588 \;lb/s"